OpenStudy (anonymous):

In a survey, 480 people, or 75%, said they attended a movie at least once a month. How many people were surveyed? A. 405 people B. 360 people C. 555 people D. 640 people

OpenStudy (anonymous):

This is a great problem to set up a proportion. We know that 480 people is 75% of some population, let's call it P. So we know that 480/P = 75/100 (because 75/100 is 75%). Therefore, we would solve 480/P=75/100, which would yield, 480*100=75P (cross multiplication). Then P=(480*100)/75. P=640. HTH!